# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 6.4 The Nested Guarded Game
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$ and let $q,m\geq 1$. Then the complement of $\operatorname{CSP}(\mathfrak{B})$ cannot be expressed by a $q$-nested width $m$ guarded query if and only if for all $k\geq 1$ there exists an unsatisfiable instance $\mathfrak{A}$ of $\operatorname{CSP}(\mathfrak{B})$ such that Duplicator has a winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$.
Proof 6.29.
For the backwards direction, let $\Pi$ be a $q$-nested guarded query of width $(m,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. By assumption, there exists an unsatisfiable instance $\mathfrak{A}$ of $\operatorname{CSP}(\mathfrak{B})$ such that Duplicator has a winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$. By the implication from 3. to 1. in Theorem 6.26, the query $\Pi$ is false on $\mathfrak{A}$. But then $\Pi$ does not express the complement of $\operatorname{CSP}(\mathfrak{B})$.
For the forwards direction, we show the contraposition. We assume that there exists a $k\geq 1$ such that for every unsatisfiable instance $\mathfrak{A}$ of $\operatorname{CSP}(\mathfrak{B})$ Duplicator has no winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$. It suffices to prove that the canonical $q$ nested guarded query $\Pi$ of width $(m,k)$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ expresses the complement of $\operatorname{CSP}(\mathfrak{B})$. Let $\mathfrak{A}$ be an instance of $\operatorname{CSP}(\mathfrak{B})$. If $\mathfrak{A}$ is satisfiable then $\Pi$ is false on $\mathfrak{A}$, because $\Pi$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$ as shown in the proof of the implication 1. implies 2. in Theorem 6.26. If $\mathfrak{A}$ is unsatisfiable then the contraposition of the implication from 2. to 3. in Theorem 6.26 implies that $\Pi$ is true on $\mathfrak{A}$. Hence, $\Pi$ expresses the complement of $\operatorname{CSP}(\mathfrak{B})$.
@@
Proof 6.31.
It is easy to see that $\Psi$ describes a class $\mathcal{C}$ of finite structures whose complement is closed under disjoint unions, and hence is the complement of a CSP. By Corollary 3.19 there exists an $\omega$-categorical structure $\mathfrak{B}$ such that the complement of $\operatorname{CSP}(\mathfrak{B})$ equals ${\mathcal{C}}$.
To prove the statement we use Corollary 6.28 for the special case $q=1$. Let $m,k\geq 1$. Let $\phi_{1}(x_{0},x_{k})$ be the formula
@@
Claim 1. $\mathfrak{A}$ satisfies $\Psi$, and hence is an unsatisfiable instance of $\operatorname{CSP}(\mathfrak{B})$. First note that $(z_{0},z_{1}),(z_{1},z_{2}),\dots,(z_{m},z_{m+1})$ are precisely the pairs in $A^{2}$ that do not satisfy $\Phi_{1}$. Then the vertices $z_{0},z_{m+1}$ play the role of $u$ and $v$: to see that $\Phi_{2}(u,v)$ does not hold in $\mathfrak{A}$, let $U_{1}\subseteq A$ be such that it satisfies the first two conjuncts of $\Phi_{2}$ for all $x,y\in A$. Then $U_{1}$ must contain $z_{0}$ by the first conjunct, and $z_{1}$ by the second conjunct, and inductively by the second conjunct it must contain all of $z_{0},z_{1},\dots,z_{m+1}$, and hence $v$. But then the last disjunct $\neg(U_{1}(x)\wedge\neg\Phi_{1}(x,v))$ is not satisfied for $x=z_{m}$.
Claim 2. Duplicator has a winning strategy ${\mathcal{H}}_{0},{\mathcal{H}}_{1}$ for the $1$-nested width $m$ guarded $k$ pebble game on $\mathfrak{A}$ and $\mathfrak{B}$: we set ${\mathcal{H}}_{0}$ to be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $\mathfrak{B}$ such that $\operatorname{dom}(h)\leq m$. Set ${\mathcal{H}}_{1}$ to be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $\mathfrak{B}$ such that if $\operatorname{dom}(h)$ contains the variable $z$ from a conjunct $\phi_{2}(y_{0},y_{k})$ of $\phi$, and $\phi_{2}^{\prime}(\bar{u})$ is the formula obtained from the rewriting of $\phi_{2}$ into prenex normal form and existentially quantifying all variables except for the ones that are in the domain of $h$, then $h(\bar{u})$ satisfies $\phi_{2}^{\prime}$ in $\mathfrak{B}$. Note that $\operatorname{dom}(h)$ cannot contain the variable $z$ for all of the conjuncts of $\phi$ of the from $\phi_{2}(y_{0},y_{k})$, and hence ${\mathcal{H}}_{1}$ is non-empty. It is straightforward to verify that ${\mathcal{H}}_{0},{\mathcal{H}}_{1}$ satisfies the three items from Definition 6.23.
@@ 6.5 Separation in GSO
Clearly, the structure $\mathfrak{B}$ is $\omega$-categorical; it is even finitely bounded homogeneous. The structure $\mathfrak{B}$ plays a prominent role in the classification of structures preserved by all permutations [11] and the complexity of quantified constraint satisfaction problems for equality constraints [10, 44]. The proof of the following proposition uses ideas from Example 2.2.
@@
Since $\operatorname{dom}(h)\leq k$ it is impossible that $r=0$ and $s=k-1$; hence, if $x_{0}$ and $x_{k}$ are from $\operatorname{dom}(h)$, then we will have $h(x_{0})\neq h(x_{k})$.
@@
Proof 6.37.
It is easy to see that $\operatorname{CSP}(\mathfrak{B})$ can be solved by a Datalog program of width $(2,3)$, because for a given $\tau$-structure $\mathfrak{A}$ it suffices to compute the smallest transitive reflexive binary relation $T$ such that for every $(x,y,z)\in R^{\mathfrak{A}}$ with $(x,y)\in T$ we have $(y,z)\in T$.
@@ 6.6 Separation in MSO
In this section we will prove that there are problems in MSO that can be expressed in Datalog, but that cannot be expressed in Nemodeq. The MSO sentence describes the complement of the CSP of a structure $\mathfrak{C}$ which is closely related to the structure $\mathfrak{B}$ from Section 6.5. Throughout this section, let $\mathfrak{C}$ be the $\{D,R\}$-structure $(C;D^{\mathfrak{C}},R^{\mathfrak{C}})$ where
@@
That is, the domain $C$ contains two sorts of elements: all natural numbers and all triples of natural numbers. [^5] For the triples inside $C$, we will use the notation ${\langle n_{1},n_{2},n_{3}\rangle}$, to better distinguish them from tuples over $C$ which we will continue to enclose in round parentheses ()($\ldots$)( … ). In words, the binary relation $D^{\mathfrak{C}}$ expresses disequality between natural numbers, whereas the quaternary relation $R^{\mathfrak{C}}$ consists of those quadruples where the first three components are natural numbers while the last component is the triple built from the first three components. Among those quadruples, $R^{\mathfrak{C}}$ contains those where the first three components are identical whenever the first two components are. An alternative formulation would be to say that $R^{\mathfrak{C}}$ contains all quadruples of the form $(x,y,z,{\langle x,y,z\rangle})$ except those where $x=y$ and $y\neq z$.
To prove that the complement of $\operatorname{CSP}(\mathfrak{C})$ is in Datalog, we show that $\mathfrak{C}$ has a primitive positive interpretation in the structure $\mathfrak{B}$ from the previous section and then use a known general transfer result for containment in Datalog.
Definition 6.40.
Let $\sigma$ and $\tau$ be finite relational signatures, and let $d,p\in{\mathbb{N}}$. A $d$-dimensional interpretation $I$ of $\tau$ in $\sigma$ consists of a $\sigma$-formula $R_{I}(\bar{x}_{1},\dots,\bar{x}_{n})$ for every $R\in\tau$ of arity $n$, where $\bar{x}_{1},\dots,\bar{x}_{n}$ are $d$-tuples of variables. If $\mathfrak{A}$ is a $\sigma$-structure, then $I(\mathfrak{A})$ is the $\tau$-structure with domain $A^{d}$ and for $R\in\tau$ the relation
$R^{I(\mathfrak{A})}:=\{(\bar{t}_{1},\dots,\bar{t}_{n})\in(A^{d})^{n}\mid\mathfrak{A}\models R_{I}(\bar{t}_{1},\dots,\bar{t}_{n})\}.$
We then say that $I$ is an interpretation of $I(\mathfrak{A})$ in $\mathfrak{A}$. If $\Theta$ is a set of formulas, then a $\Theta$-interpretation is an interpretation where all the interpreting formulas are from $\Theta$.
@@
$\mathfrak{C}$ has a 5-dimensional primitive positive interpretation in $\mathfrak{B}$.
Proof 6.42.
To define the interpretation $I$ of $\mathfrak{C}$ in $\mathfrak{B}$, let $D_{I}(x_{1},\dots,x_{5},y_{1},\dots,y_{5})$ be $x_{1}=x_{2}\wedge y_{1}=y_{2}\wedge x_{1}\neq y_{1}$. The formula $R_{I}(x_{1},\dots,x_{5},y_{1},\dots,y_{5},z_{1},\dots,z_{5},t_{1},\dots,t_{5})$ is given by
$x_{1}=x_{2}\wedge y_{1}=y_{2}\wedge z_{1}=z_{2}\wedge t_{1}\neq t_{2}$
[…]
The idea is that 5-tuples $(t_{1},\dots,t_{5})$ with $t_{1}=t_{2}$ represent elements from the sort ${\mathbb{N}}$, and 5-tuples $(t_{1},\dots,t_{5})$ with $t_{1}\neq t_{2}$ represent elements from the sort ${\mathbb{N}}^{3}$.
Corollary 6.43.
The complement of $\operatorname{CSP}(\mathfrak{C})$ is in Datalog.
Proof 6.44.
It is known that if a structure $\mathfrak{C}^{\prime}$ has a primitive positive interpretation in another structure $\mathfrak{B}^{\prime}$, and the complement of $\operatorname{CSP}(\mathfrak{B}^{\prime})$ is in Datalog, then the complement of $\operatorname{CSP}(\mathfrak{C}^{\prime})$ is in Datalog as well (proved for finite-domain structures $\mathfrak{C}$ in [37]; the same proof also works for general $\mathfrak{C}^{\prime}$, see [8, Lemma 8.3.3]). So the statement follows immediately from Lemma 6.41.
We will now argue that $\operatorname{CSP}(\mathfrak{C})$ can be expressed in MSO. We first provide a characterisation for a given $\mathfrak{A}$ being in $\operatorname{CSP}(\mathfrak{B})$.
Lemma 6.45.
A finite $\{D,R\}$-structure $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{B})$ if and only if the following conditions are satisfied:
@@
$a\mapsto\left\{\begin{array}[]{ll}f([a]_{\approx})&\text{ if }a\in A^{\prime\prime}\\ {\langle f([b]_{\approx}),f([c]_{\approx}),f([d]_{\approx})\rangle}&\text{ whenever }(b,c,d,a)\in R^{\mathfrak{A}}\\ \end{array}\right.$
@@
For the “only if” part, assume a homomorphism $h$ from $\mathfrak{A}$ to $\mathfrak{B}$. W.l.o.g. we can assume $h$ to be such that it maps any $a\in A$ not participating in any of the relations to an arbitrary element of $\mathbb{N}$. We then let $A^{\prime\prime}=h^{-1}(\mathbb{N})$ and $A^{\prime}=A\setminus A^{\prime\prime}$, which by assumption satisfy Condition (1). We define an equivalence relation $\dot{\approx}$ over $A^{\prime\prime}$ by letting $a\ \dot{\approx}\ b$ iff $h(a)=h(b)$. We observe that $\dot{\approx}$ must be disjoint from $D^{\mathfrak{A}}$ and that ${\approx}\subseteq\dot{\approx}$. Therefore $\approx$ must be disjoint from $D^{\mathfrak{A}}$, warranting Condition (2).
@@
A finite $\{D,R\}$-structure $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{B})$ if and only if it satisfies the MSO sentence $\exists X.\varphi_{\mathrm{well}}(X)$.
Proof 6.48.
For the “only if” direction, we make a formal argument along the intuitions provided above. Assume a homomorphism $h$ from $\mathfrak{A}$ to $\mathfrak{B}$. We show that the assignment $X\mapsto A^{=}$ with
@@
Validation of $\varphi_{\mathrm{triag}}(X)$ can be obtained from the fact that, for any $(a_{x},a_{y},a_{z},a_{t})\in R^{\mathfrak{A}}$ satisfying $h(a_{x})=h(a_{y})$ it must also hold that $h(a_{y})=h(a_{z})$ by virtue of $h$ being a homomorphism into $\mathfrak{B}$ – where it remains to show that $\varphi_{\mathrm{Eq}}(a_{x},a_{y},A^{=})$ indeed implies $h(a_{x})=h(a_{y})$. To see that this is true, consider the case where $Y$ is mapped to the set $A_{x}=h^{-1}(h(a_{x}))$. Then the implication’s consequence indeed ensures $h(a_{x})=h(a_{y})$, provided that the premise is found to be true. The premise’s first conjunct holds by construction, whence it remains to show that $\varphi_{\mathrm{propag}}(A_{x})$ also holds true, which – in view of our chosen assignment for $X$ – again follows from the fact that $h$ is a homomorphism into $\mathfrak{B}$
Finally, validation of $\varphi_{\mathrm{compat}}(X)$ is established by recalling that $\varphi_{\mathrm{Eq}}(a_{x},a_{y},A^{=})$ implies $h(a_{x})=h(a_{y})$ (see above) and therefore $(h(a_{x}),h(a_{y}))\not\in D^{\mathfrak{B}}$ by definition.
For the “if” direction, assume that some assignment $X\mapsto A_{X}\subseteq A$ validates $\varphi_{\mathrm{well}}(X)$ for $\mathfrak{A}$. We will use Lemma 6.45 to argue that $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{B})$. Let $A^{\prime\prime}=\{a_{4}\mid(a_{1},a_{2},a_{3},a_{4})\in R^{\mathfrak{A}}\}$ and let $A^{\prime}=A\setminus A^{\prime\prime}$. Then, by virtue of $\varphi_{\mathrm{sort}}$, the partition $(A^{\prime},A^{\prime\prime})$ satisfies Condition (1) of Lemma 6.45. Satisfaction of the Condition (2) would be an immediate consequence of $\varphi_{\mathrm{compat}}(X)$ if one could establish that the binary relation $Eq_{A_{X}}\coloneqq\{(a_{1},a_{2})\mid\varphi_{\mathrm{Eq}}(a_{1},a_{2},A_{X})\}$ contains $\approx$. This is achieved by arguing that $Eq_{A_{X}}$ is an equivalence relation and satisfies the two subitems in Lemma 6.45 defining $\approx$. The former is a result of the structure of $\varphi_{\mathrm{Eq}}$ as an undirected reachability check. The latter is enforced through $\varphi_{\mathrm{propag}}$ and $\varphi_{\mathrm{triag}}$ which together ensure the two subconditions (2a) and (2b).

### v3

@@ 5.5 The Nested Guarded Game
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$ and let $q,m\geq 1$. Then the complement of $\operatorname{CSP}(\mathfrak{B})$ cannot be expressed by a $q$-nested width $m$ guarded query if and only if for all $k\geq 1$ there exists a finite $\tau$-structure $\mathfrak{A}\notin\operatorname{CSP}(\mathfrak{B})$ such that Duplicator has a winning strategy for the $q$-nested width $m$ guarded $k$-pebble game on $(\mathfrak{A},\mathfrak{B})$.
Proof 5.30.
For the backwards direction, let $\Pi$ be a $q$-nested guarded query of width $(m,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. By assumption, there exists $\mathfrak{A}\notin\operatorname{CSP}(\mathfrak{B})$ such that Duplicator has a winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$. By the implication from 3. to 1. in Theorem 5.27, the query $\Pi$ is false on $\mathfrak{A}$. But then $\Pi$ does not express the complement of $\operatorname{CSP}(\mathfrak{B})$.
For the forwards direction, we show the contraposition. We assume that there exists a $k\geq 1$ such that for every $\mathfrak{A}\notin\operatorname{CSP}(\mathfrak{B})$ Duplicator has no winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$. It suffices to prove that the canonical $q$ nested guarded query $\Pi$ of width $(m,k)$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ expresses the complement of $\operatorname{CSP}(\mathfrak{B})$. Let $\mathfrak{A}$ be an instance of $\operatorname{CSP}(\mathfrak{B})$. If $\mathfrak{A}$ is satisfiable then $\Pi$ is false on $\mathfrak{A}$, because $\Pi$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$ as shown in the proof of the implication 1. implies 2. in Theorem 5.27. If $\mathfrak{A}$ is unsatisfiable then the contraposition of the implication from 2. to 3. in Theorem 5.27 implies that $\Pi$ is true on $\mathfrak{A}$. Hence, $\Pi$ expresses the complement of $\operatorname{CSP}(\mathfrak{B})$.
To illustrate the power of our game characterisation of GQ, we show that there are problems in the intersection of Datalog and MSO that are inexpressible in GQ.
@@
Proof 5.32.
It is easy to see that $\Psi$ describes a class $\mathcal{C}$ of finite structures whose complement is closed under disjoint unions, and hence is the complement of a CSP by Remark 1.2. By Corollary 3.24 there exists an $\omega$-categorical structure $\mathfrak{B}$ such that the complement of $\operatorname{CSP}(\mathfrak{B})$ equals $\mathcal{C}$}.
To prove the inexpressibility in GQ, we use Corollary 5.29 for the special case $q=1$. Let $m,k\geq 1$. Let $\phi_{1}(x_{0},x_{k})$ be the formula
@@
Claim 1. $\mathfrak{A}$ satisfies $\Psi$, and hence is an unsatisfiable instance of $\operatorname{CSP}(\mathfrak{B})$. First note that $(z_{0},z_{1}),(z_{1},z_{2}),\dots,(z_{m},z_{m+1})$ are precisely the pairs in $A^{2}$ that do not satisfy the formula $\Phi_{1}$ from definition of $\Psi$ in Example 5.11. Then the vertices $z_{0},z_{m+1}$ play the role of $u$ and $v$: to see that $\Phi_{2}(u,v)$ does not hold in $\mathfrak{A}$, let $U_{1}\subseteq A$ be such that it satisfies the first two conjuncts of $\Phi_{2}$ for all $x,y\in A$. Then $U_{1}$ must contain $z_{0}$ by the first conjunct, and $z_{1}$ by the second conjunct, and inductively by the second conjunct it must contain all of $z_{0},z_{1},\dots,z_{m+1}$, and hence $v$. But then the last disjunct $\neg(U_{1}(x)\wedge\neg\Phi_{1}(x,v))$ is not satisfied for $x=z_{m}$.
Claim 2. Duplicator has a winning strategy ${\mathcal{H}}_{0},{\mathcal{H}}_{1}$ for the $1$-nested width $m$ guarded $k$ pebble game on $\mathfrak{A}$ and $\mathfrak{B}$: we set ${\mathcal{H}}_{0}$ to be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $\mathfrak{B}$ such that $\operatorname{dom}(h)\leq m$. Set ${\mathcal{H}}_{1}$ to be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $\mathfrak{B}$ such that if $\operatorname{dom}(h)$ contains the variable $z$ from a conjunct $\phi_{2}(y_{0},y_{k})$ of $\phi$, and $\phi_{2}^{\prime}(\bar{u})$ is the formula obtained from the rewriting of $\phi_{2}$ into prenex normal form and existentially quantifying all variables except for the ones that are in the domain of $h$, then $h(\bar{u})$ satisfies $\phi_{2}^{\prime}$ in $\mathfrak{B}$. Note that $\operatorname{dom}(h)$ cannot contain the variable $z$ for all of the conjuncts of $\phi$ of the form $\phi_{2}(y_{0},y_{k})$, and hence ${\mathcal{H}}_{1}$ is non-empty. It is straightforward to verify that ${\mathcal{H}}_{0},{\mathcal{H}}_{1}$ satisfies the three items from Definition 5.23.
@@ 5.6 Separation in GSO
Clearly, the structure $\mathfrak{B}$ is $\omega$-categorical. The structure $\mathfrak{B}$ plays a prominent role in the classification of structures preserved by all permutations [13] and the complexity of quantified constraint satisfaction problems for equality constraints [12, 45]. The proof of the following proposition uses ideas from Example 2.2.
@@
Since $\operatorname{dom}(h)\leq k$ it is impossible that $r=0$ and s=k1s=k-1; hence, if $x_{0}$ and $x_{k}$ are from $\operatorname{dom}(h)$, then we will have $h(x_{0})\neq h(x_{k})$.
@@
Proof 5.38.
It is easy to see that the complement of $\operatorname{CSP}(\mathfrak{B})$ can be defined by a Datalog program of width $(2,3)$, because for a given $\tau$-structure $\mathfrak{A}$ it suffices to
  • compute the smallest reflexive, symmetric, and transitive binary relation $T$ on all elements participating in $R$ such that for every $(x,y,z)\in R^{\mathfrak{A}}$ with $(x,y)\in T$ we have $(y,z)\in T$:
    $T(y,z)$ $\;{:}{-}\;T(x,y),R(x,y,z)$ $T(y,x)$ $\;{:}{-}\;T(x,y)$ $T(x,z)$ $\;{:}{-}\;T(x,y),T(y,z)$ $T(x,x)$ $\;{:}{-}\;R(x,y,z)$ $T(y,y)$ $\;{:}{-}\;R(x,y,z)$ $T(z,z)$ $\;{:}{-}\;R(x,y,z),$
  • and then to check whether $T$ and $D$ contain a common pair of elements or $D$ has a loop:
    $\qquad\ \ {\sf goal}\;{:}{-}\;T(x,y),D(x,y)\qquad\qquad\ {\sf goal}\;{:}{-}\;D(x,x).$
@@ 5.7 Separation in MSO
In this section, we will prove that there are problems in MSO that can be expressed in Datalog, but that cannot be expressed in Nemodeq. The MSO sentence describes the complement of the CSP of a structure $\mathfrak{C}$ which is closely related to the structure $\mathfrak{B}$ from Section 5.6. Throughout this section, let $\mathfrak{C}$ be the $\{D,R\}$-structure $(C;D^{\mathfrak{C}},R^{\mathfrak{C}})$ where
@@
That is, the domain $C$ contains two sorts of elements: all natural numbers and all triples of natural numbers. [^9] For the triples inside $C$, we will use the notation $\langle n_{1},n_{2},n_{3}\rangle$}, to better distinguish them from tuples over $C$ which we will continue to enclose in round parentheses $(\ldots)$. In words, the binary relation $D^{\mathfrak{C}}$ expresses disequality between natural numbers, whereas the quaternary relation $R^{\mathfrak{C}}$ consists of quadruples where the first three components are natural numbers while the last component is the triple built from the first three components. Among all such quadruples, $R^{\mathfrak{C}}$ contains those where the first three components are identical whenever the first two components are. An alternative formulation would be to say that $R^{\mathfrak{C}}$ contains all quadruples of the form $(x,y,z,{\langle x,y,z\rangle})$ except those where $x=y$ and $y\neq z$.
Proposition 5.41.
The complement of $\operatorname{CSP}(\mathfrak{C})$ is in Datalog.
Proof 5.42.
The idea of the proof is analoguous to that of Proposition 5.37. Consider the following Datalog program $\Pi$
$T(y,z)$ $\;{:}{-}\;T(x,y),R(x,y,z,t)$ $T(y,x)$ $\;{:}{-}\;T(x,y)$ $T(x,z)$ $\;{:}{-}\;T(x,y),T(y,z)$ ${\sf goal}$ $\;{:}{-}\;T(x,y),D(x,y)$ ${\sf goal}$ $\;{:}{-}\;T(t,t),R(x,y,z,t)$ $T(x,x^{\prime})$ $\;{:}{-}\;R(x,y,z,t),R(x^{\prime},y^{\prime},z^{\prime},t)$ $T(y,y^{\prime})$ $\;{:}{-}\;R(x,y,z,t),R(x^{\prime},y^{\prime},z^{\prime},t)$ $T(z,z^{\prime})$ $\;{:}{-}\;R(x,y,z,t),R(x^{\prime},y^{\prime},z^{\prime},t)$ $T(x,x)$ $\;{:}{-}\;D(x,y)$ $T(y,y)$ $\;{:}{-}\;D(x,y)$
Now assume $\mathfrak{A}$ is such that $\Pi$ does not derive $\sf goal$} and let $\mathfrak{A}^{\prime}=\Pi(\mathfrak{A})$. Let
$A^{*}=\{a_{1},a_{2},a_{3}\in A\mid(a_{1},a_{2},a_{3},a_{4})\in R^{\mathfrak{A}}\}\cup\{a_{1},a_{2}\in A\mid(a_{1},a_{2})\in D^{\mathfrak{A}}\}.$
Thanks to $\mathfrak{A}^{\prime}=\Pi(\mathfrak{A})$ and $\Pi(\mathfrak{A})$ not deriving $\sf goal$}, we observe the following:
  1. 1.
    $T^{\mathfrak{A}^{\prime}}$ is an equivalence relation on $A^{*}$ and is disjoint from $D^{\mathfrak{A}}$,
  2. 2.
    $A^{*}$ is disjoint from $\{a_{4}\in A\mid(a_{1},a_{2},a_{3},a_{4})\in R^{\mathfrak{A}}\}$,
  3. 3.
    for any $(a_{1},a_{2},a_{3},a)\in R^{\mathfrak{A}}$ and $(a^{\prime}_{1},a^{\prime}_{2},a^{\prime}_{3},a)\in R^{\mathfrak{A}}$, we have that $a_{i}$ and $a^{\prime}_{i}$ are in the same $T^{\mathfrak{A}^{\prime}}$-equivalence class for any $i\in\{1,2,3\}$.
We now define a homomorphism $h$ from $\mathfrak{A}$ to $\mathfrak{C}$. Pick an injective function $h^{*}\colon{A^{*}}_{\!\!/T^{\mathfrak{A}^{\prime}}}\to\mathbb{N}$ mapping $T^{\mathfrak{A}^{\prime}}$-equivalence classes to natural numbers. For every $a\in A^{*}$, we let $h(a)=h^{*}([a]_{T^{\mathfrak{A}^{\prime}}})$ and for every $a\in A\setminus A^{*}$, we let $h(a)\in\mathbb{N}^{3}$ be such that, whenever $(a_{1},a_{2},a_{3},a)\in R^{\mathfrak{A}}$ for some $a_{1},a_{2},a_{3}\in A^{*}$ then $h(a)=\langle h(a_{1}),h(a_{2}),h(a_{3})\rangle$ (thanks to item 3 above, this choice is always possible). It can now be readily checked that $h$ is a homomorphism from $\mathfrak{A}$ to $\mathfrak{C}$.
We will now argue that $\operatorname{CSP}(\mathfrak{C})$ can be expressed in MSO. We first provide a characterisation for a given $\mathfrak{A}$ being in $\operatorname{CSP}(\mathfrak{C})$.
@@
A finite $\{D,R\}$-structure $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{C})$ if and only if the following conditions are satisfied:
@@
[…]
@@
[…]
@@
A finite $\{D,R\}$-structure $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{C})$ if and only if it satisfies the MSO sentence $\exists X.\varphi_{\mathrm{well}}(X)$.
Proof 5.46.
For the “only if” direction, we make a formal argument along the intuitions provided above. Assume a homomorphism $h$ from $\mathfrak{A}$ to $\mathfrak{C}$. We show that the assignment $X\mapsto A^{=}$ with
@@
Validation of $\varphi_{\mathrm{triag}}(X)$ can be obtained from the fact that, for any $(a_{x},a_{y},a_{z},a_{t})\in R^{\mathfrak{A}}$ satisfying $h(a_{x})=h(a_{y})$ it must also hold that $h(a_{y})=h(a_{z})$ by virtue of $h$ being a homomorphism into $\mathfrak{C}$ – where it remains to show that $\varphi_{\mathrm{Eq}}(a_{x},a_{y},A^{=})$ indeed implies $h(a_{x})=h(a_{y})$. To see that this is true, consider the case where $Y$ is mapped to the set $A_{x}=h^{-1}(h(a_{x}))$. Then the implication’s consequence indeed ensures $h(a_{x})=h(a_{y})$, provided that the premise is found to be true. The premise’s first conjunct holds by construction, whence it remains to show that $\varphi_{\mathrm{propag}}(A_{x})$ also holds true, which – in view of our chosen assignment for $X$ – again follows from the fact that $h$ is a homomorphism into $\mathfrak{C}$ .
Finally, validation of $\varphi_{\mathrm{compat}}(X)$ is established by recalling that $\varphi_{\mathrm{Eq}}(a_{x},a_{y},A^{=})$ implies $h(a_{x})=h(a_{y})$ (see above) and therefore $(h(a_{x}),h(a_{y}))\not\in D^{\mathfrak{C}}$ by definition.
For the “if” direction, assume that some assignment $X\mapsto A_{X}\subseteq A$ validates $\varphi_{\mathrm{well}}(X)$ for $\mathfrak{A}$. We will use Lemma˜5.43 to argue that $\mathfrak{A}$ is in $\operatorname{CSP}(\mathfrak{C})$. Let $A^{\prime\prime}=\{a_{4}\mid(a_{1},a_{2},a_{3},a_{4})\in R^{\mathfrak{A}}\}$ and let $A^{\prime}=A\setminus A^{\prime\prime}$. Then, by virtue of $\varphi_{\mathrm{sort}}$, the partition $(A^{\prime},A^{\prime\prime})$ satisfies Condition (1) of Lemma˜5.43. Satisfaction of the Condition (2) would be an immediate consequence of $\varphi_{\mathrm{compat}}(X)$ if one could establish that the binary relation $Eq_{A_{X}}\coloneqq\{(a_{1},a_{2})\mid\varphi_{\mathrm{Eq}}(a_{1},a_{2},A_{X})\}$ contains $\approx$. This is achieved by arguing that $Eq_{A_{X}}$ is an equivalence relation and satisfies the two subitems in Lemma˜5.43 defining $\approx$. The former is a result of the structure of $\varphi_{\mathrm{Eq}}$ as an undirected reachability check. The latter is enforced through $\varphi_{\mathrm{propag}}$ and $\varphi_{\mathrm{triag}}$ which together ensure the two subconditions (2a) and (2b).